OpenStudy (anonymous):

Find the Limit as x approaches infinity. Sqrt(x^2-3x)-x? I know the answer is 1.5 I just don't know how to reach that...the right way

OpenStudy (anonymous):

multiply by \[\frac{\sqrt{x^2-3x}+x}{\sqrt{x^2-3x}+x}\]

OpenStudy (anonymous):

haha, oh wow, thank you :)

OpenStudy (anonymous):

you get \[\frac{x^2-3x-x^2}{\sqrt{x^2-3x}+x}\] or \[\frac{-3x}{\sqrt{x^2-3x}+x}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i actually get \(-1.5\) though